# Higher equations of motion at level 2 in Liouville CFT

### v1

@@ 2.1.2. Semigroups and Poisson operator
By Kahane’s theorey of Gaussian multiplicative chaos (GMC) [Kah85], we can define the random measure
$\mathrm{d}M_{\gamma}(z)=\underset{\epsilon\to 0}{\lim}\,\epsilon^{\frac{\gamma^{2}}{2}}e^{\gamma X_{\epsilon}(z)},$
where $X_{\epsilon}=X\star\rho_{\epsilon}$ is a convolution regularisation of the field (with $\rho$ a bump function and $\rho_{\epsilon}(z)=\epsilon^{-2}\rho(\frac{z}{\epsilon})$).
There are two important one-parameter semigroups of operators on $\mathcal{H}$, with respective generators denoted by $\mathbf{H}^{0}$ and $\mathbf{H}$. Both generators are positive, essentially self-adjoint, unbounded operators. They are called the free field and Liouville semigroups respectively. The operator $\mathbf{H}$ defines a positive definite quadratic form with domain $\mathcal{D}(\mathcal{Q})\subset\mathcal{H}$. The continuous dual of $\mathcal{D}(\mathcal{Q})$ is denoted by $\mathcal{D}^{\prime}(\mathcal{Q})$. Explicitly, the semigroups are defined by the probabilistic formula:
$e^{-t\mathbf{H}^{0}}F(c+\varphi)=e^{-\frac{Q^{2}}{2}t}\mathbb{E}_{\varphi}\left[F(c+X(e^{-t}\cdot\,))\right]$
$e^{-t\mathbf{H}}F(c+\varphi)=e^{-\frac{Q^{2}}{2}t}\mathbb{E}_{\varphi}\left[F(c+X(e^{-t}\cdot\,)e^{-\mu e^{\gamma c}\int_{\mathbb{A}_{t}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma Q}}}\right],$
where $\mathbb{A}_{t}=\{e^{-t}<|z|<1\}$, and $\mathbb{E}_{\varphi}$ means the conditional expectation with respect to $\varphi$ (i.e. $c+\varphi$ is fixed and we integrate over $X_{\mathbb{D}}$).
Informally, the Poisson operator is a map sending free field eigenstates to Liouville eigenstates. The eigenstates of $\mathbf{H}^{0}$ are easily described: for each $p\in\mathbb{R}$, $e^{ipc}\mathcal{F}_{N}$ is an eigenspace with eigenvalue $2\Delta_{Q+ip}+N$. The diagonalisation of $\mathbf{H}$ is much harder and is the cornerstone of the conformal bootstrap theorem [GKRV23]. The Poisson operator is constructed using the long term asymptotics of the Liouville semigroup as follows. Let $\nu,\widetilde{\nu}\in\mathcal{T}$. For all $\alpha$ in a certain neighbourhood of $\infty$- ∞, the limit
$\mathcal{P}_{\alpha}(\mathcal{Q}_{\alpha,\nu,\widetilde{\nu}}):=\underset{t\to\infty}{\lim}\,e^{t(2\Delta_{\alpha}+|\nu|+|\widetilde{\nu}|)}e^{-t\mathbf{H}}\left(\mathcal{Q}_{\alpha,\nu,\widetilde{\nu}}e^{(\alpha-Q)c}\right)$
exists in a weighted space $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. In this region, we define $\Phi_{\alpha}:=\mathcal{P}_{\alpha}\circ\Phi_{\alpha}^{0}$, and we have the explicit expression
@@
where $\Psi_{\alpha,\nu,\widetilde{\nu}}$ are generalised eigenstates of the Liouville Hamiltonian [GKRV23], with eigenvalue $2\Delta_{\alpha}+|\nu|+|\widetilde{\nu}|$. It is known that $\alpha\mapsto\Psi_{\alpha,\nu,\widetilde{\nu}}$ extends analytically to the whole plane and satisfies the reflection formula $\Psi_{2Q-\alpha,\nu,\widetilde{\nu}}=R(\alpha)\Psi_{\alpha,\nu,\widetilde{\nu}}$, with $R(\alpha)$ the reflection coefficient [${}^{+}$23]. Thus, for each $\nu,\widetilde{\nu}\in\mathcal{T}$, the map $\alpha\mapsto\Phi_{\alpha}(\pi_{\mathbf{k},\widetilde{\mathbf{k}}})$ extends analytically to the whole plane. Since $\alpha\mapsto\Phi_{\alpha}^{0}$ is analytic with zeros located on the Kac table, it follows that $\alpha\mapsto\mathcal{P}_{\alpha}$ extends meromorphically to the whole plane, with possible poles on the Kac table.
For $\alpha$ in a complex neighbourhood of $(-\infty,Q)$, the highest-weight state $\Psi_{\alpha}$ has a probabilistic representation
$\Psi_{\alpha}=e^{(\alpha-Q)c}\mathbb{E}_{\varphi}\left[e^{-\mu e^{\gamma c}\int_{\mathbb{D}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}}\right].$
Following [BW23], for $r\in\mathbb{N}^{*}$ and non-coinciding points $w_{1},...,w_{r}\in\mathbb{D}\setminus\{0\}$, we introduce the notation
@@
[…]
@@ 2.2. Expression of the singular state
This can be proved by Gaussian integration by parts, which we view as a differential version of the Girsanov transform. We follow the method of [BW23, Section 3.5]. Let $\varepsilon\in\mathbb{C}$ and consider the martingale
@@
with initial value $\mathcal{E}_{\varepsilon}(0)=e^{2\mathrm{Re}(\varepsilon\varphi_{2})}$. Observe that
@@
The effect of reweighting the measure by $e^{-2\mathrm{Re}(\varepsilon\varphi_{2})}\mathcal{E}_{\varepsilon}(t)$ is to shift the Dirichlet free field as
@@
Moreover, by the Girsanov transform, we have for all $F\in\mathcal{C}$:
$\mathbb{E}\left[e^{2\mathrm{Re}(\varepsilon\varphi_{2})-\frac{|\varepsilon|^{2}}{4}}F(\varphi)\right]=\mathbb{E}\left[F\left(\varphi+\frac{1}{2}\mathrm{Re}(\varepsilon e_{-2})\right)\right].$
@@
Hence, the total shift of the field $X=X_{\mathbb{D}}+P\varphi$ is $X(z)\mapsto X(z)+\mathrm{Re}(\varepsilon z^{-2})$. For all $t>0$, we then get
$e^{t(2\Delta_{\alpha}+4)}\mathbb{E}\left[e^{-t\mathbf{H}}\left((4|\varphi_{2}|^{2}-1)e^{(\alpha-Q)c}\right)F(\varphi)\right]$
@@
$\quad=4e^{(\alpha-Q)c}\frac{\partial^{2}}{\partial\varepsilon\partial\bar{\varepsilon}}_{|\varepsilon=0}e^{-\frac{|\varepsilon|^{2}}{4}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}\mathcal{E}_{\varepsilon}(t)e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}F(\varphi)\right]$
@@
$\quad=4e^{(\alpha-Q)c}\frac{\partial^{2}}{\partial\varepsilon\partial\bar{\varepsilon}}_{|\varepsilon=0}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{-\mu e^{\gamma c}\int_{\mathbb{A}_{t}}e^{\frac{\gamma}{2}\mathrm{Re}(\varepsilon z^{-2})}\mathrm{d}M_{\gamma}(z)}F(\varphi+\frac{1}{2}\mathrm{Re}(\varepsilon e_{-2}))\right]$
@@
$\quad=-\frac{\mu\gamma^{2}}{4}e^{(\alpha+\gamma-Q)c}\int_{\mathbb{A}_{t}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{\gamma X(w)-\frac{\gamma^{2}}{2}\mathbb{E}[X(w)^{2}]}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}F(\varphi)\right]\frac{|\mathrm{d}w|^{2}}{|w|^{4}}$
@@
$\qquad+\frac{\mu^{2}\gamma^{2}}{4}e^{(\alpha+2\gamma-Q)c}\int_{\mathbb{A}_{t}}\int_{\mathbb{A}_{t}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{\gamma X(w_{1})-\frac{\gamma^{2}}{2}\mathbb{E}[X(w_{1})^{2}]}e^{\gamma X(w_{2})-\frac{\gamma^{2}}{2}\mathbb{E}[X(w_{2})]^{2}}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}F(\varphi)\right]\frac{|\mathrm{d}w_{1}|^{2}}{\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}^{2}}$
@@
$\qquad-\frac{\mu\gamma}{4}e^{(\alpha+\gamma-Q)c}\int_{\mathbb{A}_{t}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{\gamma X(w)-\frac{\gamma^{2}}{2}\mathbb{E}[X(w)^{2}]}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}\partial_{2}F(\varphi)\right]\frac{|\mathrm{d}w|^{2}}{w^{2}}$
@@
$\qquad-\frac{\mu\gamma}{4}e^{(\alpha+\gamma-Q)c}\int_{\mathbb{A}_{t}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{\gamma X(w)-\frac{\gamma^{2}}{2}\mathbb{E}[X(w)^{2}]}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}\partial_{-2}F(\varphi)\right]\frac{|\mathrm{d}w|^{2}}{\bar{w}^{2}}$
@@
$\qquad+\frac{1}{4}e^{(\alpha-Q)c}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}\partial_{2}\partial_{-2}F(\varphi)\right]$
$\underset{t\to\infty}{\to}-\frac{\mu\gamma^{2}}{4}\mathbb{E}[\mathcal{I}_{1,(2),(2)}(\alpha)F(\varphi)]+\frac{\mu^{2}\gamma^{2}}{4}\mathbb{E}[\mathcal{I}_{2,(2,0),(0,2)}(\alpha)F(\varphi)]+\mathbb{E}[\mathcal{R}_{\alpha}F],$
@@
where $\mathcal{R}_{\alpha}$ denotes the sum of the last three lines. By [BW23, Proposition 3.2], $\mathcal{R}_{\alpha}$ is analytic for $\alpha$ in a complex neighbourhood of $(-\infty,0)$, which concludes the proof by density of $\mathcal{C}$ in $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. ∎
@@ 2.3. First pole of $\mathcal{P}_{\alpha}$ and the $(1,2)$-HEM
Let $F\in\mathcal{C}$. By Proposition 2.6 applied to $r=1$, we have in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$
@@
for some $\xi>0$, uniformly in a neighbourhood of $\alpha_{1,2}$. Using this estimate, we have (recall $\alpha_{1,2}+\gamma=\alpha_{-1,2}$)
$\mathbb{E}[\mathcal{I}_{1,(2),(2)}(\alpha)F]$ $=\mathbb{E}[\Psi_{\alpha_{-1,2}}F]\int_{\mathbb{D}}|w|^{-\gamma\alpha}\frac{|\mathrm{d}w|^{2}}{|w|^{4}}+\int_{\mathbb{D}}\mathbb{E}[(\Psi_{\alpha}(w)-|w|^{-\gamma\alpha}\Psi_{\alpha_{-1,2}})F]\frac{|\mathrm{d}w|^{2}}{|w|^{4}}$ (2.2)
$=-\frac{2\pi}{\gamma(\alpha-\alpha_{1,2})}\mathbb{E}[\Psi_{\alpha_{-1,2}}F]+O_{\alpha\to\alpha_{1,2}}(1),$
with the $O(1)$ coming from the integrability of $|w|^{-4-\gamma\alpha+\xi}$ in the neighbourhood of $\alpha_{1,2}$.
This shows that $\mathbb{E}[\mathcal{I}_{1,(2),(2)}(\alpha)F]$ is meromorphic in a neighbourhood of $\alpha_{1,2}$ with a simple pole there, and $\underset{\alpha=\alpha_{1,2}}{\mathrm{Res}}\,\mathcal{I}_{1,(2),(2)}(\alpha)=-\frac{2\pi}{\gamma}\mathbb{E}[\Psi_{\alpha_{-1,2}}]$. Combining with Proposition 2.2, we get $\underset{\alpha=\alpha_{1,2}}{\mathrm{Res}}\,\mathcal{P}_{\alpha}(4|\varphi_{2}|^{2}-1)=\pi\mu\frac{\gamma}{2}\Psi_{\alpha_{-1,2}}$ in $\mathcal{C}^{\prime}$.
Moreover, we know that $\mathcal{P}_{\alpha}(4|\varphi_{2}|^{2}-1)$ is meromorphic with values in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$, so the residue exists in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. Moreover, it is clear from (2.2) that $(\alpha-\alpha_{1,2})\mathcal{I}_{1,(2),(2)}(\alpha)$ is bounded in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$ in a neighbourhood of $\alpha_{1,2}$, so that the residue holds in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. ∎
@@
$\frac{\partial}{\partial\alpha}_{|\alpha=\alpha_{1,2}}\mathcal{P}_{\alpha}(\mathbf{S}^{0}_{\alpha}\widetilde{\mathbf{S}}^{0}_{\alpha}\mathds{1})$ $=\underset{\alpha\to\alpha_{1,2}}{\lim}\,\frac{1}{\alpha-\alpha_{1,2}}\mathcal{P}_{\alpha}(\mathbf{S}^{0}_{\alpha}\widetilde{\mathbf{S}}^{0}_{\alpha}\mathds{1})$
$=\underset{\alpha\to\alpha_{1,2}}{\lim}\,(\alpha-\alpha_{1,2})\alpha^{2}(\alpha-\alpha_{2,1})^{2}\mathcal{P}_{\alpha}(4|\varphi_{2}|^{2}-1)$

### v2

@@ 2.1.2. Semigroups and Poisson operator
By Kahane’s theory of Gaussian multiplicative chaos (GMC) [KAH85] and variants [DS11, BER17, RV14], we can define the random measure
$\mathrm{d}M_{\gamma}(z)=\underset{\epsilon\to 0}{\lim}\,\epsilon^{\frac{\gamma^{2}}{2}}e^{\gamma X_{\epsilon}(z)}|\mathrm{d}z|^{2},$
where $X_{\epsilon}$ is a regularisation of the field at scale $\epsilon>0$. For concreteness, we will use the circle average regularisation [BP24, Section 1.12], but the limiting measure is unique among a large choice of regularisations [BER17, Theorem 1.1].
There are two important one-parameter semigroups of operators on $\mathcal{H}$, with respective generators denoted by $\mathbf{H}^{0}$ and $\mathbf{H}$. Both generators are positive, self-adjoint, unbounded operators. They are called the free field and Liouville semigroups respectively. The operator $\mathbf{H}$ defines a positive definite quadratic form with domain $\mathcal{D}(\mathcal{Q})\subset\mathcal{H}$. The continuous dual of $\mathcal{D}(\mathcal{Q})$ is denoted by $\mathcal{D}^{\prime}(\mathcal{Q})$. Explicitly, the semigroups are defined by the probabilistic formula, for all $F\in\mathcal{H}$:
$e^{-t\mathbf{H}^{0}}F(c+\varphi)=e^{-\frac{Q^{2}}{2}t}\mathbb{E}_{\varphi}\left[F(c+X(e^{-t}\cdot\,))\right]$ (2.4)
$e^{-t\mathbf{H}}F(c+\varphi)=e^{-\frac{Q^{2}}{2}t}\mathbb{E}_{\varphi}\left[F(c+X(e^{-t}\cdot\,))e^{-\mu e^{\gamma c}\int_{\mathbb{A}_{t}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma Q}}}\right],$
where $\mathbb{A}_{t}=\{e^{-t}<|z|<1\}$, and $\mathbb{E}_{\varphi}$ means the conditional expectation with respect to $\varphi$ (i.e. $c+\varphi$ is fixed and we integrate over $X_{\mathbb{D}}$). See [GKR+24a, Sections 4.1 & 4.3] for more details on the free field semigroup, and [GKR+24a, Proposition 5.1] for the Liouville semigroup.
Informally, the Poisson operator is a map sending free field (generalised) eigenstates to Liouville (generalised) eigenstates. The eigenstates of $\mathbf{H}^{0}$ are easily described: for each $p\in\mathbb{R}$, $e^{ipc}\mathcal{F}_{N}$ is an eigenspace with eigenvalue $2\Delta_{Q+ip}+N$. The diagonalisation of $\mathbf{H}$ is much harder and is the cornerstone of the conformal bootstrap theorem [GKR+24a]. The Poisson operator is constructed using the long time asymptotics of the Liouville semigroup as follows. Let $\nu,\widetilde{\nu}\in\mathcal{T}$. For all $\alpha\in\mathbb{C}$ with $\mathrm{Re}(\alpha)$ sufficiently small, the limit
[…]
exists in a weighted space $e^{-\beta c}\mathcal{D}(\mathcal{Q})$ with $\beta>Q-\mathrm{Re}(\alpha)$ [BGK+24a, (3.26)]. In this region, we define $\Phi_{\alpha}:=\mathcal{P}_{\alpha}\circ\Phi_{\alpha}^{0}$ (recall (2.3) for the definition of $\Phi_{\alpha}^{0}$), and we have the explicit expression
@@
[…]
For $\alpha\in(-\infty,Q)$, the highest-weight state $\Psi_{\alpha}$ has a probabilistic representation [BGK+24a, (1.7)] [^2] In the sequel, we will consider states with additional parameters, so we will write $\Psi_{\alpha}$ for $\Psi_{\alpha}(c,\varphi)$ to lighten the notation.
$\Psi_{\alpha}(c,\varphi)=e^{(\alpha-Q)c}\mathbb{E}_{\varphi}\left[e^{-\mu e^{\gamma c}\int_{\mathbb{D}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}}\right],$
and it admits an analytic continuation in $\alpha$ [BGK+24a, Theorem 1.2]. Following [BW23, (3.2)], for $r\in\mathbb{Z}_{>0}$ and non-coinciding points $w_{1},...,w_{r}\in\mathbb{D}\setminus\{0\}$, we introduce the notation
@@
which is the primary field $\Psi_{\alpha}$ with additional $\gamma$-insertions at $w_{1},...,w_{r}$. By the Cameron–Martin shift applied to the exponential prefactors, the limit can be expressed as a Laplace transform of Gaussian multiplicative chaos with singularities at $0,w_{1},...,w_{r}$ [BW23, (3.3)]. In fact, we will only focus on the case $r\;\leqslant\;2$. Given $\boldsymbol{s}=(s_{1},...,s_{r}),\widetilde{\boldsymbol{s}}=(\widetilde{s}_{1},...,\widetilde{s}_{r})\in\mathbb{N}^{r}$, we define, if it exists
@@ 2.2. Expression of the singular state
This can be proved by Gaussian integration by parts, which we view as a differential version of the Girsanov transform, following the method of [BW23, Lemma 3.11]. Under the free field dynamics generated by $\mathbf{H}^{0}$, the time evolution of the field can be represented as
$c_{t}+\varphi_{t}(e^{i\theta})=c+P\varphi(e^{-t+i\theta})+X_{\mathbb{D}}(e^{-t+i\theta}),$
where $X_{\mathbb{D}}$ is an independent Dirichlet GFF in $\mathbb{D}$. Moreover, all the modes of the field are independent, with $c_{t}=c+B_{t}$ evolving a standard Brownian motion started from $c$, and $\mathrm{Re}(\varphi_{n}(t))$, $\mathrm{Im}(\varphi_{n}(t))$ evolving as Ornstein–Uhlenbeck processes (where $\varphi_{n}(t)$ is defined through the expansion of $\varphi_{t}$ in Fourier modes: $\varphi_{t}(e^{i\theta})=2\mathrm{Re}(\sum_{n=1}^{\infty}\varphi_{n}(t)e^{ni\theta})$). See [GKR+24a, Sections 4.1 & 4.3] for details.
For each $\varepsilon\in\mathbb{C}$, we can then consider the exponential martingale:
@@
with initial value $\mathcal{E}_{\varepsilon}(0)=e^{2\mathrm{Re}(\varepsilon\varphi_{2})}$. This martingale was already introduced in the proof of [BW23, Lemma 3.11]. Observe that
@@
By Girsanov’s theorem, the effect of reweighting the measure by $e^{-2\mathrm{Re}(\varepsilon\varphi_{2})}\mathcal{E}_{\varepsilon}(t)$ is to shift the Dirichlet free field as
@@
Moreover, by the Cameron–Martin formula, we have for all $F\in\mathcal{C}$ (recall the standard basis introduced in (2.1)):
$\mathbb{E}\left[e^{2\mathrm{Re}(\varepsilon\varphi_{2})-\frac{|\varepsilon|^{2}}{4}}F(c,\varphi)\right]=\mathbb{E}\left[F\left(c,\varphi+\frac{1}{2}\mathrm{Re}(\varepsilon e_{-2})\right)\right].$
@@
Hence, the total shift of the field $X=X_{\mathbb{D}}+P\varphi$ is $X(z)\mapsto X(z)+\mathrm{Re}(\varepsilon z^{-2})$. For all $t>0$, the Feynman–Kac formula of [GKR+24a, Proposition 5.1] and Girsanov’s theorem give
$e^{t(2\Delta_{\alpha}+4)}\mathbb{E}\left[e^{-t\mathbf{H}}\left((4|\varphi_{2}|^{2}-1)e^{(\alpha-Q)c}\right)F(c,\varphi)\right]$
@@
$\quad=4e^{(\alpha-Q)c}\frac{\partial^{2}}{\partial\varepsilon\partial\bar{\varepsilon}}_{|\varepsilon=0}e^{-\frac{|\varepsilon|^{2}}{4}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}\mathcal{E}_{\varepsilon}(t)e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}F(c,\varphi)\right]$
@@
$\quad=4e^{(\alpha-Q)c}\frac{\partial^{2}}{\partial\varepsilon\partial\bar{\varepsilon}}_{|\varepsilon=0}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{-\mu e^{\gamma c}\int_{\mathbb{A}_{t}}e^{\frac{\gamma}{2}\mathrm{Re}(\varepsilon z^{-2})}\mathrm{d}M_{\gamma}(z)}F(c,\varphi+\frac{1}{2}\mathrm{Re}(\varepsilon e_{-2}))\right]$
@@
$\quad=-\frac{\mu\gamma^{2}}{4}e^{(\alpha+\gamma-Q)c}\int_{\mathbb{A}_{t}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{\gamma X(w)-\frac{\gamma^{2}}{2}\mathbb{E}[X(w)^{2}]}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}F(c,\varphi)\right]\frac{|\mathrm{d}w|^{2}}{|w|^{4}}$
@@
$\qquad+\frac{\mu^{2}\gamma^{2}}{4}e^{(\alpha+2\gamma-Q)c}\int_{\mathbb{A}_{t}}\int_{\mathbb{A}_{t}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{\gamma X(w_{1})-\frac{\gamma^{2}}{2}\mathbb{E}[X(w_{1})^{2}]}e^{\gamma X(w_{2})-\frac{\gamma^{2}}{2}\mathbb{E}[X(w_{2})]^{2}}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}F(c,\varphi)\right]\frac{|\mathrm{d}w_{1}|^{2}}{\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}^{2}}$
@@
$\qquad-\frac{\mu\gamma}{4}e^{(\alpha+\gamma-Q)c}\int_{\mathbb{A}_{t}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{\gamma X(w)-\frac{\gamma^{2}}{2}\mathbb{E}[X(w)^{2}]}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}\partial_{2}F(c,\varphi)\right]\frac{|\mathrm{d}w|^{2}}{w^{2}}$
@@
$\qquad-\frac{\mu\gamma}{4}e^{(\alpha+\gamma-Q)c}\int_{\mathbb{A}_{t}}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{\gamma X(w)-\frac{\gamma^{2}}{2}\mathbb{E}[X(w)^{2}]}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}\partial_{-2}F(c,\varphi)\right]\frac{|\mathrm{d}w|^{2}}{\bar{w}^{2}}$
@@
$\qquad+\frac{1}{4}e^{(\alpha-Q)c}\mathbb{E}\left[e^{\alpha B_{t}-\frac{\alpha^{2}}{2}t}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t})}\partial_{2}\partial_{-2}F(c,\varphi)\right]$
$\underset{t\to\infty}{\to}-\frac{\mu\gamma^{2}}{4}\mathbb{E}[\mathcal{I}_{1,(2),(2)}(\alpha)F]+\frac{\mu^{2}\gamma^{2}}{4}\mathbb{E}[\mathcal{I}_{2,(2,0),(0,2)}(\alpha)F]+\mathbb{E}[\mathcal{R}_{\alpha}F].$
@@
For $\alpha$ very negative, the above convergence is readily seen using the Cameron–Martin shift and the representations of the states in [BW23, (3.3)]. By [BW23, Proposition 3.2], $\mathcal{R}_{\alpha}$ is analytic for $\alpha$ in a complex neighbourhood of $(-\infty,0)$, which concludes the proof by density of $\mathcal{C}$ in $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. ∎
@@ 2.3. First pole of $\mathcal{P}_{\alpha}$ and the $(1,2)$-HEM
Let $F\in\mathcal{C}$. By Proposition 2.6 applied to $r=1$, we have in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$:
@@
for some $\xi>0$, uniformly in a neighbourhood of $\alpha_{1,2}$. Using this estimate, we have
$\mathbb{E}[\mathcal{I}_{1,(2),(2)}(\alpha)F]$ $=\mathbb{E}[\Psi_{\alpha+\gamma}F]\int_{\mathbb{D}}|w|^{-\gamma\alpha}\frac{|\mathrm{d}w|^{2}}{|w|^{4}}+\int_{\mathbb{D}}\mathbb{E}[(\Psi_{\alpha}(w)-|w|^{-\gamma\alpha}\Psi_{\alpha+\gamma})F]\frac{|\mathrm{d}w|^{2}}{|w|^{4}}$ (2.6)
$=-\frac{2\pi}{\gamma(\alpha-\alpha_{1,2})}\mathbb{E}[\Psi_{\alpha+\gamma}F]+O_{\alpha\to\alpha_{1,2}}(1),$
with the $O(1)$ coming from the integrability of $|w|^{-4-\gamma\alpha+\xi}$ in $\mathbb{D}$, for $\alpha$ in a neighbourhood of $\alpha_{1,2}$.
This shows that $\mathbb{E}[\mathcal{I}_{1,(2),(2)}(\alpha)F]$ is meromorphic in a neighbourhood of $\alpha_{1,2}$ with a simple pole there, so that $\underset{\alpha=\alpha_{1,2}}{\mathrm{Res}}\,\mathcal{I}_{1,(2),(2)}(\alpha)=-\frac{2\pi}{\gamma}\Psi_{\alpha_{-1,2}}$ in $\mathcal{C}^{\prime}$ (recall $\alpha_{-1,2}=\alpha+\gamma$). Combining with Proposition 2.2, we get $\underset{\alpha=\alpha_{1,2}}{\mathrm{Res}}\,\mathcal{P}_{\alpha}(4|\varphi_{2}|^{2}-1)=\pi\mu\frac{\gamma}{2}\Psi_{\alpha_{-1,2}}$ in $\mathcal{C}^{\prime}$.
Moreover, we have explained in Section 2.1 that $\mathcal{P}_{\alpha}(4|\varphi_{2}|^{2}-1)$ is meromorphic with values in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$, so the residue exists in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. Now, the function $(\alpha-\alpha_{1,2})\mathcal{I}_{1,(2),(2)}(\alpha)$ is analytic in a neighbourhood of $\alpha_{1,2}$ so that Cauchy’s formula gives $(\alpha-\alpha_{1,2})\mathcal{I}_{1,(2),(2)}(\alpha)=\frac{1}{2i\pi}\oint\mathcal{I}_{1,(2),(2)}(\alpha^{\prime})\frac{\alpha^{\prime}-\alpha_{1,2}}{\alpha^{\prime}-\alpha}\mathrm{d}\alpha^{\prime}$ in $\mathcal{C}^{\prime}$ in a neighbourhood of $\alpha=\alpha_{1,2}$, where the contour is a small loop surrounding $\alpha_{1,2}$. Now, the right-hand-side is an integral of elements in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$ and avoids the singularity at $\alpha_{1,2}$, so it converges to $\frac{1}{2i\pi}\oint\mathcal{I}_{1,(2),(2)}(\alpha)\mathrm{d}\alpha=\underset{\alpha=\alpha_{1,2}}{\mathrm{Res}}\mathcal{I}_{1,(2),(2)}(\alpha)$ in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. ∎
@@
$\frac{\partial}{\partial\alpha}_{|\alpha=\alpha_{1,2}}\mathcal{P}_{\alpha}(\mathbf{S}^{0}_{\alpha}\widetilde{\mathbf{S}}^{0}_{\alpha}\mathds{1})=\underset{\alpha\to\alpha_{1,2}}{\lim}\,\frac{1}{\alpha-\alpha_{1,2}}\mathcal{P}_{\alpha}(\mathbf{S}^{0}_{\alpha}\widetilde{\mathbf{S}}^{0}_{\alpha}\mathds{1})$ $=\underset{\alpha\to\alpha_{1,2}}{\lim}\,(\alpha-\alpha_{1,2})\alpha^{2}(\alpha-\alpha_{2,1})^{2}\mathcal{P}_{\alpha}(4|\varphi_{2}|^{2}-1)$